OpenStudy (anonymous):

A 29-foot ladder is leaning against a building. The base of the ladder is 20 feet from the side of the building. How high does the ladder reach along the side of the building?

OpenStudy (lgbasallote):

 use pythagorean theorem to solve for x

OpenStudy (lgbasallote):

\[\huge x = \sqrt{29^2 - 20^2}\]

OpenStudy (anonymous):

what is the pyethagorean theorem before you plugged numbers into it?

OpenStudy (anonymous):

That S^2 = r^2 + r^r and that r^2 = s^2 - r^2 I a right triangle where s = the side accross the 90 degree corner and the r = any of the other two sides of the triangle.